OpenStudy (anonymous):

does anybody know about extended set operations and indexed families of sets

OpenStudy (anonymous):

\[\displaystyle B -\left ( \bigcup _{\alpha \in \Delta} A_{\alpha} \right ) = \bigcup_{\alpha \in \Delta} \left (B - A_{\alpha} \right ) \]